# CorDA: Context-Oriented Decomposition Adaptation of Large Language Models for Task-Aware Parameter-Efficient Fine-tuning

### v2

@@ 3.1 Preliminaries on Low-Rank Adaptation
Figure 1: An overall illustration of our proposed method. We perform singular value decomposition oriented by the covariance matrix to aggregate task context into the principle components (up), which are frozen for maintaining world knowledge (down left) or utilized to initialize the learnable adapter for better fine-tuning performance (down right). The dark-colored adapter refers to the components with the largest $r$ singular values, while the light one is composed of the smallest $r$ components.
@@ 3.3 Mode 1: Knowledge-Preserved Adaptation
where $B\in\mathbb{R}^{d_{out}\times r}$ and $A\in\mathbb{R}^{r\times d_{in}}$ are the initialized matrices in the learnable adapter, $BA=\sum_{i=R-r+1}^{R}\sigma_{i}\mathbf{u}_{i}\mathbf{\hat{v}}_{i}^{T}$ corresponds to the last $r$ components in Eq. (3), $\sqrt{\Sigma}_{[-r:]}$ is a diagonal matrix with the squared root of the smallest $r$ singular values on the diagonal, and $W^{\prime}$ corresponding to the first $R-r$ components in Eq. (3) is frozen during fine-tuning. We have $W^{\prime}$ as the difference between $W$ and $BA$ instead of summing the first $R-r$ components to avoid the numerical error between $\hat{W}$ and $W$ introduced by the decomposition and inversion operations. After fine-tuning, the learned matrices $B^{*}$ and $A^{*}$ can be merged into $W^{\prime}$ as $W^{*}=W^{\prime}+B^{*}A^{*}$. This mode is featured by preserving world knowledge as the knowledge retrieving ability captured by the principle $R-r$ components is frozen. It is also more effective than a zero-initialized adapter in learning new abilities as verified by our experiments.
@@ 4.1 Analysis of the Ability to Capture Context
We conduct an experiment to demonstrate the ability of our proposed decomposition method in Sec. 3.2 to capture context in its principle components. We use different methods including the Plain SVD, ASVD [73], and our Context-Oriented SVD (CO-SVD), to perform full decomposition of the LLaMA-2-7B pre-trained weights in all layers, and then discard the smallest $r$ singular values and their corresponding left and right singular vectors to reconstruct the weights for testing.
As shown in Figure 2, as the number of discarded ranks increases, the performances with Plain SVD on both Wikitext-2 [48] and PTB [46] are getting worse steeply. ASVD considers data context using activation absolute mean values, and helps to relieve the deterioration compared to Plain SVD. However, when discarding more than 256 ranks, the Perplexity also diverges sharply. In contrast, our method is able to maintain a stable performance very close to the original pre-trained weights even when the smallest 1024 components are discarded. The result indicates that our method is proficient with aggregating context from limited samples (256 Wiki or PTB samples in this example) into the principle components. It also evidences the potential of our method to maintain important knowledge when freezing the principle components and to learn new abilities when adapting these components. The detailed numbers of the experiment in Figure 2 are listed in Table 6 (Appendix B), where we also test the effect of sample number and dataset choice when collecting the covariance matrices.
@@ 4.2 Knowledge-Preserved Adaptation Results
4.3 Instruction-Previewed Adaptation Results
@@ 4.4 Discussions
Ablations. We ablate the data choice used to produce covariance matrices and the adapter building manner in Table 4. The first row corresponds to the implementation of PiSSA that uses the largest $r$ singular values and their singular vectors as the initialized adapter by plain SVD. If we use the smallest $r$ components by plain SVD to build adapters, there is no apparent improvement in world knowledge benchmarks. This implies that the plain SVD cannot precisely capture world knowledge related ability into the principle components and thus freezing them does not help to mitigate knowledge forgetting. When our context-oriented decomposition (CO-SVD) is adopted with Wikitext-2, which is not closely correlated with question answering, the performance on world knowledge is much improved. When the context collected by the covariance matrix is from question answering data, i.e., TriviaQA or NQ open, the world knowledge performance is further improved by a significant margin, and the average score is also enhanced as a result. Therefore, data context is important to orientate the decomposition process such that the characteristics of the ability concerned can be better aggregated into the principle components for maintaining or adapting. As shown in Table 5, similar to TriviaQA and NQ open in the knowledge-preserved adaptation, collecting context from different data sources belonging to the same category, namely WizardLM-Evol-Instruct and Alpaca, also results in close performance in the instruction-previewed adaptation. It is noteworthy that CorDA with Alpaca on MTBench (5.06) is still the highest among the compared baselines in Table 2.
@@ 5 Conclusion
In this paper, we propose a new parameter-efficient fine-tuning method, named context-oriented decomposition adaptation (CorDA). It performs singular value decomposition for pre-trained weights oriented by the covariance matrix that captures the context of the task concerned, and aggregates the context into the principle components for maintaining or adapting. Accordingly, our method is able to support two implementation modes, the knowledge-preserved adaptation to mitigate world knowledge forgetting and the instruction-previewed adaptation for better fine-tuning performance. In experiments, our knowledge-preserved adaptation not only achieves better fine-tuning performance than LoRA, but also maintains the world knowledge well, leading to the best average scores on three fine-tuning tasks. Our instruction-previewed adaptation is able to further enhance the fine-tuning performance, surpassing the state-of-the-art parameter-efficient fine-tuning methods DoRA and PiSSA.
@@
This work was also supported in part by the National Natural Science Foundation of China under Grant 62376069, in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, and in part by Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012027.
@@ Appendix B Appendix: More Results
Table 6 lists the detailed numbers and more results of the experiment in Figure 2. It is shown that the number of sampled data only has a very limited impact. When the smallest 1024 ranks are discarded, using 32 samples is slightly worse than 256 samples in both Wikitext-2 and PTB. It implies that a small number of samples is enough to capture context into the principle components. Besides, collecting samples from the same dataset as the one used to test is able to attain a better performance after discarding a large number of ranks. For example, when discarding the smallest 1024 ranks, CO-SVD (with 256 Wiki samples) is better than CO-SVD (with 256 PTB samples) on Wikitext-2 (6.35 v.s. 8.69), and CO-SVD (with 256 PTB samples) is better than CO-SVD (with 256 Wiki samples) on PTB (22.28 v.s. 29.69). This also reveals that precisely capturing the data context in our decomposition is crucial for better maintaining the task characteristics into the principle components, and explains why our method is superior to the PEFT methods without considering data context.

### v3

@@ 3.1 Preliminaries on Low-Rank Adaptation
Figure 1: An overall illustration of our proposed method. We perform singular value decomposition oriented by the covariance matrix to aggregate task context into the principal components (up), which are frozen for maintaining world knowledge (down left) or utilized to initialize the learnable adapter for better fine-tuning performance (down right). The dark-colored adapter refers to the components with the largest $r$ singular values, while the light one is composed of the smallest $r$ components.
@@ 3.3 Mode 1: Knowledge-Preserved Adaptation
where $B\in\mathbb{R}^{d_{out}\times r}$ and $A\in\mathbb{R}^{r\times d_{in}}$ are the initialized matrices in the learnable adapter, $BA=\sum_{i=R-r+1}^{R}\sigma_{i}\mathbf{u}_{i}\mathbf{\hat{v}}_{i}^{T}$ corresponds to the last $r$ components in Eq. (3), $\sqrt{\Sigma}_{[-r:]}$ is a diagonal matrix with the squared root of the smallest $r$ singular values on the diagonal, and $W^{\prime}$ corresponding to the first $R-r$ components in Eq. (3) is frozen during fine-tuning. We have $W^{\prime}$ as the difference between $W$ and $BA$ instead of summing the first $R-r$ components to avoid the numerical error between $\hat{W}$ and $W$ introduced by the decomposition and inversion operations. After fine-tuning, the learned matrices $B^{*}$ and $A^{*}$ can be merged into $W^{\prime}$ as $W^{*}=W^{\prime}+B^{*}A^{*}$. This mode is featured by preserving world knowledge as the knowledge retrieving ability captured by the principal $R-r$ components is frozen. It is also more effective than a zero-initialized adapter in learning new abilities as verified by our experiments.
@@ 4.1 Analysis of the Ability to Capture Context
We conduct an experiment to demonstrate the ability of our proposed decomposition method in Sec. 3.2 to capture context in its principal components. We use different methods including the Plain SVD, ASVD [73], and our Context-Oriented SVD (CO-SVD), to perform full decomposition of the LLaMA-2-7B pre-trained weights in all layers, and then discard the smallest $r$ singular values and their corresponding left and right singular vectors to reconstruct the weights for testing.
As shown in Figure 2, as the number of discarded ranks increases, the performances with Plain SVD on both Wikitext-2 [48] and PTB [46] are getting worse steeply. ASVD considers data context using activation absolute mean values, and helps to relieve the deterioration compared to Plain SVD. However, when discarding more than 256 ranks, the Perplexity also diverges sharply. In contrast, our method is able to maintain a stable performance very close to the original pre-trained weights even when the smallest 1024 components are discarded. The result indicates that our method is proficient with aggregating context from limited samples (256 Wiki or PTB samples in this example) into the principal components. It also evidences the potential of our method to maintain important knowledge when freezing the principal components and to learn new abilities when adapting these components. The detailed numbers of the experiment in Figure 2 are listed in Table 8 (Appendix B), where we also test the effect of sample number and dataset choice when collecting the covariance matrices.
@@ 4.2 Knowledge-Preserved Adaptation Results
The results of CorDA in knowledge-preserved adaptation mode (fine-tuning on Math while preserving NQ open) using other LLM scales and architectures are shown in Table 2.
Table 2: The results of CorDA in knowledge-preserved adaptation mode (KPM) fine-tuning on Math while preserving NQ open with various LLM scales and architectures.
Model Method NQ open GSM8k Math Avg
LLaMA-2-13B Pre-trained 23.63 - - -
LoRA [21] 16.26 57.24 8.92 27.47
CorDA (KPM) 19.86 59.29 9.62 29.59
LLaMA-3-8B Pre-trained 13.41 - - -
LoRA [21] 8.75 72.33 24.04 35.04
CorDA (KPM) 9.61 74.68 25.34 36.54
Gemma-2-9B Pre-trained 12.85 - - -
LoRA [21] 9.28 83.47 42.30 45.02
CorDA (KPM) 10.17 84.08 42.64 45.63
@@
4.3 Instruction-Previewed Adaptation Results
The results of CorDA in instruction-previewed adaptation mode (fine-tuning on Math) using other LLM scales and architectures are shown in Table 4.
@@ 4.3 Instruction-Previewed Adaptation Results
Table 4: The results of CorDA in instruction-previewed adaptation mode (IPM) fine-tuning on Math with various LLM scales and architectures.
Model Method GSM8k Math
LLaMA-2-13B LoRA [21] 57.24 8.92
PiSSA [47] 60.88 11.08
CorDA (IPM) 62.47 11.54
Gemma-2-9B LoRA [21] 83.47 42.30
PiSSA [47] 84.23 43.52
CorDA (IPM) 84.45 43.88
@@ 4.4 Discussions
Ablations. We ablate the data choice used to produce covariance matrices and the adapter building manner in Table 6. The first row corresponds to the implementation of PiSSA that uses the largest $r$ singular values and their singular vectors as the initialized adapter by plain SVD. If we use the smallest $r$ components by plain SVD to build adapters, there is no apparent improvement in world knowledge benchmarks. This implies that the plain SVD cannot precisely capture world knowledge related ability into the principal components and thus freezing them does not help to mitigate knowledge forgetting. When our context-oriented decomposition (CO-SVD) is adopted with Wikitext-2, which is not closely correlated with question answering, the performance on world knowledge is much improved. When the context collected by the covariance matrix is from question answering data, i.e., TriviaQA or NQ open, the world knowledge performance is further improved by a significant margin, and the average score is also enhanced as a result. Therefore, data context is important to orientate the decomposition process such that the characteristics of the ability concerned can be better aggregated into the principal components for maintaining or adapting. As shown in Table 7, similar to TriviaQA and NQ open in the knowledge-preserved adaptation, collecting context from different data sources belonging to the same category, namely WizardLM-Evol-Instruct and Alpaca, also results in close performance in the instruction-previewed adaptation. It is noteworthy that CorDA with Alpaca on MTBench (5.06) is still the highest among the compared baselines in Table 3.
@@ 5 Conclusion
In this paper, we propose a new parameter-efficient fine-tuning method, named context-oriented decomposition adaptation (CorDA). It performs singular value decomposition for pre-trained weights oriented by the covariance matrix that captures the context of the task concerned, and aggregates the context into the principal components for maintaining or adapting. Accordingly, our method is able to support two implementation modes, the knowledge-preserved adaptation to mitigate world knowledge forgetting and the instruction-previewed adaptation for better fine-tuning performance. In experiments, our knowledge-preserved adaptation not only achieves better fine-tuning performance than LoRA, but also maintains the world knowledge well, leading to the best average scores on three fine-tuning tasks. Our instruction-previewed adaptation is able to further enhance the fine-tuning performance, surpassing the state-of-the-art parameter-efficient fine-tuning methods DoRA and PiSSA.
@@
This work was also supported in part by the National Natural Science Foundation of China under Grant 62376069, in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, in part by Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012027, and in part by the Shenzhen Science and Technology Program (Grant No. ZDSYS20230626091203008).
@@ Appendix B Appendix: More Results
Table 8 lists the detailed numbers and more results of the experiment in Figure 2. It is shown that the number of sampled data only has a very limited impact. When the smallest 1024 ranks are discarded, using 32 samples is slightly worse than 256 samples in both Wikitext-2 and PTB. It implies that a small number of samples is enough to capture context into the principal components. Besides, collecting samples from the same dataset as the one used to test is able to attain a better performance after discarding a large number of ranks. For example, when discarding the smallest 1024 ranks, CO-SVD (with 256 Wiki samples) is better than CO-SVD (with 256 PTB samples) on Wikitext-2 (6.35 v.s. 8.69), and CO-SVD (with 256 PTB samples) is better than CO-SVD (with 256 Wiki samples) on PTB (22.28 v.s. 29.69). This also reveals that precisely capturing the data context in our decomposition is crucial for better maintaining the task characteristics into the principal components, and explains why our method is superior to the PEFT methods without considering data context.